OpenStudy (anonymous):

Find Zeros of f(x) = −5 cos(2x +π/3)?

OpenStudy (anonymous):

I should equal to 0, but still can't get the answer

OpenStudy (anonymous):

\[{\pi n \over2} - {5\pi \over 12}\]

OpenStudy (anonymous):

It says the answer is wrong(

OpenStudy (anonymous):

sorry, that's the best I could do

OpenStudy (anonymous):

x= π/12 and x= π/3

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

π/12 is one of the many many solutions... (the one where N=1)

OpenStudy (anonymous):

Unfortunately, it's incorrect(